# New Applications of Ergodic Theory to Sets of Differences

### v1

@@ 5.3 Tempered sequences in Locally Finite Groups
Suppose that $G$ is a countably infinite, locally finite group. Enumerate its elements $g_{1},g_{2},g_{3},\ldots$ and for all $n$ let $G_{n}$ be the (finite) subgroup generated by the elements $g_{1},\ldots,g_{n}$. Then $(G_{n})_{n=1}^{\infty}$ is a tempered Følner sequence in $G$. The next two examples give instances of this idea being applied to locally finite noncommutative groups.
Example 5.7. Let $S$ denote the group of all finite permutations of $\mathbb{N}$: $S=\{\sigma:\mathbb{N}\rightarrow\mathbb{N}\quad:\quad\sigma(n)=n\text{ for allbut finitely many }n\}$. For $n\in\mathbb{N}$ the symmetric group $S_{n}$ is a subgroup of $S$ and in fact $S=\bigcup_{n=1}^{\infty}S_{n}$. If $\sigma\in S$ then $\sigma\in S_{m}$ for some $m$, so for $n>m$ we have $|S_{n}\cap\sigma S_{n}|=|S_{n}|$. It follows that $|S_{n}\cap\sigma S_{n}|/|S_{n}|\rightarrow 1$, hence $(S_{n})_{n=1}^{\infty}$ is a Følner sequence. For $n$, since $S_{n}$ is a group, we have $S_{n-1}^{-1}S_{n}\subseteq S_{n}$, so $(S_{n})_{n=1}^{\infty}$ is tempered.
@@
For $n\in\mathbb{N}$ define
$F_{n}=\left\{\begin{pmatrix}1&f&h\\ 0&1&g\\ 0&0&1\end{pmatrix}:\ \deg(f),\deg(g)\leq n\text{ and }\deg(h)\leq 2n\right\}$

### v2

@@ 5.4 Tempered sequences in Locally Finite Groups
Suppose that $G$ is a countably infinite, locally finite group. Let $g_{1},g_{2},g_{3},\ldots$ be an enumeration of its elements and for all $n\in\mathbb{N},$ let $G_{n}$ be the (finite) subgroup generated by the elements $g_{1},\ldots,g_{n}$. Then $(G_{n})_{n=1}^{\infty}$ is a tempered Følner sequence in $G$. The next two examples give instances of this idea being applied to locally finite non-commutative groups.
Example 5.8. Let $S$ denote the group of all finite permutations of $\mathbb{N}$:
$S=\{\sigma:\mathbb{N}\rightarrow\mathbb{N},\,\sigma(n)=n\text{ for all but finitely many }n\}.$
For $n\in\mathbb{N}$ the symmetric group $S_{n}$ is a subgroup of $S$ and, in fact, $S=\bigcup_{n=1}^{\infty}S_{n}$. If $\sigma\in S$, then $\sigma\in S_{m}$ for some $m$, so for $n>m$ we have $|S_{n}\cap\sigma S_{n}|=|S_{n}|$. It follows that $|S_{n}\cap\sigma S_{n}|/|S_{n}|\rightarrow 1$, hence $(S_{n})_{n=1}^{\infty}$ is a Følner sequence. For any $n\in\mathbb{N}$, since $S_{n}$ is a group, we have $S_{n-1}^{-1}S_{n}\subseteq S_{n}$, so $(S_{n})_{n=1}^{\infty}$ is tempered.
@@
For $n\in\mathbb{N}$, define
$F_{n}=\left\{\begin{pmatrix}1&f&h\\ 0&1&g\\ 0&0&1\end{pmatrix}:\ \deg(f),\deg(g)\leq n\text{ and }\deg(h)\leq 2n\right\}.$